# ULSR-GS: Ultra Large-scale Surface Reconstruction Gaussian Splatting with Multi-View Geometric Consistency

### v1

@@ 7.1 Additional implementation details
7.2 Additional partition comparisons
Novel view synthesis. Table 6 quantitatively compares different partitioning methods, specifically our point-based partitioning approach and the image-based method employed by Vast Gaussian [19]. The results indicate that the Vast Gaussian partitioning strategy is more suitable for rendering tasks. This is because our method prioritizes the reconstruction of the scene’s primary structure by excluding the distant sparse SfM points during sub-region partitioning. Figure 9 provides a qualitative comparison to further illustrate this difference. When training ULSR-GS using different partitioning strategies, our partition method shows a trade-off: while it loses performance in reconstructing the background, it preserves significantly more detail in the reconstruction of the primary scene components.
Figure 10: Qualitative comparison of partition strategies.
Figure 11: Qualitative comparison of partition strategies.
Figure 12: Qualitative comparison of partition strategies at scene edges. We train our method on our partition strategy and Vast Gaussian’s [19] in $8\times 8$ cells. From top to bottom: example scene edge, example GT images at the scene edge, ULSR-GS on our partition strategy, and ULSR-GS on Vast Gaussian’s [19] partition strategy.
Mesh extraction. Since UrbanScene 3D [21] does not provide a reference ground truth point cloud, we perform a qualitative analysis. And Figures 10 and 11 qualitatively demonstrate comparisons on both five-directional oblique photogrammetry datasets and close-range oblique photogrammetry datasets. Our partitioning strategy effectively identifies and selects the most critical images for the reconstruction regions, significantly enhancing the detail and accuracy of the extracted meshes.
Figure 12 demonstrates a qualitative analysis of partition strategies at the edges of reconstructed scenes. Comparison was performed on the campus scene from the UrbanScene3D [21] dataset. Our partition strategy demonstrates the ability to provide sufficient training images even at the periphery of the scene. In cases of uneven camera distribution (particularly at the periphery of the scene), the image-based partitioning may lead to insufficient coverage of edge regions.
8 Detailed Additional Densification Implementations
8.1 Implementations
Our depth projection process is integrated into the training pipeline after the default densification phase of the vanilla 2DGS [11]. The densification process employs geometric consistency checks with thresholds set to geo_pixel_thres = 1 and geo_depth_thres = 0.001 to ensure robust depth projection. We set the proportional constant of the window size to $k$=0.6 in our experiments.
The densification begins at the 20,000th iteration to ensure that accurate rendered depths are available prior to projection. During each densification iteration, we perform local scale re-initialization within the projection area to better adapt to the scene’s depth variations. To balance computational efficiency and model performance, we reduce the interval of opacity culling operations to 2,000 iterations. This adjustment prevents over-fitting by limiting the excessive accumulation of projected points while maintaining sufficient pruning to minimize computational overhead.
9 Addition Results
9.1 Mesh evaluation process
Since GauU-scene [38] not provide a standard evaluation code, we follow the Tanks and Tample [14] evaluation pipeline to quantitatively assess reconstructed meshes against GT LiDAR point clouds:
  1. 1.
  2. 2.
    We uniformly sample number of points from the reconstructed mesh and GT points to 10 million.
  3. 3.
    We utilize KDTreeFlann to compute precision, recall, and F-score using predefined distance thresholds ($\tau=0.025,0.01,0.005$).
Figure 13: Qualitative comparisons between two depth calculations. Although using median depth in our method can improve the geometry of the building, a large number of holes will appear in the plane area.
9.2 Additional mesh results
$\tau=0.01$ $\tau=0.005$
Metrics Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$ Pre.$\uparrow$ Rec.$\uparrow$ F1$\uparrow$
Neuralangelo * * * * * *
SuGaR 0.246 0.243 0.244 0.212 0.209 0.210
2DGS-60K 0.403 0.432 0.416 0.235 0.267 0.250
GOF-60K 0.429 0.451 0.439 0.262 0.273 0.267
PGSR-60K 0.417 0.429 0.422 0.249 0.253 0.250
Ours (Median depth) 0.432 0.461 0.447 0.262 0.271 0.266
Ours 0.429 0.467 0.451 0.261 0.280 0.269
Table 7: Quantitative results on GauU-Scene dataset [38]. We report precision, recall, and F-1 score at the threshold $\tau=0.01,0.005$. The best, second best, and third best results are highlighted in red, orange, and yellow respectively. ”-” denotes no results due to GPU out of memory. ” *” denotes no result due to no convergence.
To further evaluate the reconstruction accuracy, we conducted experiments on the GauU-Scene dataset [38] using stricter thresholds of $\tau=0.01$ and $\tau=0.005$ as presented in Table 7. Our method consistently outperforms existing approaches across both thresholds. At $\tau=0.01$ , our method achieves the highest precision and recall, resulting in the best F-score.
Notably, as the evaluation threshold $\tau$ becomes more stringent (from 0.01 to 0.005), the performance metrics of all models tend to converge, indicating increased difficulty in achieving high precision and recall. Despite this convergence, our method maintains a leading edge, achieving the highest recall and F1 score at $\tau=0.005$.
9.3 Depth comparison
Table 7 shows the comparison results of mesh extraction using two different methods. Using median depth can further improve the precision of the mesh, but the overall index has decreased, which leads to a decrease in the overall F1 score. Figure 13 further reveals this result, which can be seen that although using median depth can visually improve the fineness of the mesh, it results in depth calculation errors for some smooth objects, such as the ground, resulting in a large number of holes in the mesh.
Figure 14: Full scene comparison. We show the full scene comparison with 2DGS [11], GOF[47], and PGSR[3]. From left to right are the LFLS, SZIIT, and SMBU scenes of the GauU-Scene [38] dataset.
9.4 Additional qualitative results
Examples of the partitioning mesh results can be found in Figure 19. A full scene comparison on the GauU-Scene [38] dataset among our method, 2DGS [11], GOF[47], and PGSR[3] can be found in Figure 14. Detailed qualitative comparisons between 2DGS [11], GOF[47], PGSR[3], and SuGaR [10] are provided in Figure 15, 16, 17, and 18, respectively.
10 Detailed Conclusion and Limitations
Partition. Our approach addresses key limitations of image-based partitioning methods [19, 22, 13] by ensuring finer granularity in mesh boundaries and prioritizing the reconstruction of key scene regions. This strategy enables us to extract meshes individually rather than merge sub-regions and extract the full scene. Through both quantitative and qualitative evaluations, we demonstrated significant improvements in reconstruction accuracy and detail preservation, particularly in close-range path planning photogrammetry datasets.
However, our method exhibits certain limitations. Specifically, the emphasis on prioritizing the reconstruction of key areas within the scene inadvertently diminishes the fidelity of distant background regions. While this approach is advantageous for tasks like mesh extraction that focus on scene details, it results in suboptimal performance for rendering tasks that demand comprehensive scene coverage, including distant and sparsely observed areas. Future work can focus on addressing these limitations by extending the partition strategy to achieve a balance between reconstruction fidelity in critical regions and the preservation of distant background elements.
Depth Projection Densification. Our depth projection densification strategy primarily targets improving the quality of well-reconstructed regions, enhancing their rendering and reconstruction fidelity. However, it does not address under-reconstructed areas effectively. Additionally, the reliance on multi-view geometric consistency constraints can lead to instability in reflective regions, such as water surfaces and glass. These limitations highlight the need for future work to develop more robust densification and constraints strategies that not only address under-reconstructed areas but also improve the handling of challenging reflective surfaces.
Figure 15: Qualitative comparisons between ours and 2DGS [11].
Figure 16: Qualitative comparisons between ours and GOF [47].
Figure 17: Qualitative comparisons between ours and PGSR [3].
Figure 18: Qualitative comparisons between ours and SuGaR [10].
Figure 19: Example of partition mesh results. Top to bottom: SMBU, SZIIT from GauU-Scene [38], and Hospital from UrbanScene 3D [21]

### v2

@@ 5.6.3 Impact of Partition Strategies
6 Discussion
6.1 Real World Application: Integration with USV Bathymetry model
We now discuss the ability of integrating the bathymetry digital elevation model (BDEM) obtained from single-beam echo sounder (SBES) with the oblique photogrammetric models produced by ULSR-GS to generate a real urban surface model. As shown in Figure 17 (a), the experimental site for this study is located at Moon Bay (Scene 2), situated along the southern shore of Dushu Lake in Suzhou, China. Moon Bay is a scenic and ecologically rich area, but it is also surrounded by a dense cluster of urban buildings, which presents significant challenges for UAV-based photogrammetric reconstruction and the subsequent data fusion with USV-derived bathymetric models.
Figure 17: Experimental Location. (a) Bird-eye-view of the experimental location Moon Bay, Dushu Lake, Suzhou. (b) USV and its SBES mathymetry footprints. (c) UAV and its oblique photogrammetry footprints.
As shown in Figure 17 (b), the USV followed predefined paths across the water body, recording depth measurements at each point along its trajectory. The WGS-84 world coordinates were recorded with each measurement, providing a geospatial reference for each bathymetry point. Simultaneously, a UAV was deployed to capture oblique imagery of the surrounding landscape. Each image was georeferenced using the UAV’s embedded GPS, recording the WGS-84 world coordinates at the time of capture. The initial data consists of bathymetry measurements at specific points. The bathymetric data collected by the USV is first processed using Geographic Information System (GIS) software, where the raw measurements are converted into a DEM through Kriging interpolation. The high-resolution DEM is converted into a 3D mesh using standard Triangulated Irregular Network (TIN) algorithms. As shown in Figure 18, the ULSR-GS-generated mesh and the UAV-derived mesh are then aligned and merged. The alignment process ensures that both meshes are represented in the same global coordinate system (WGS-84 Mercator projection), allowing for seamless integration.
Figure 18: Zoomed-in detail of the fusion USV and UAV models. Left: Mesh; Right Normal.
6.2 Strength and Limitations
Our proposed ULSR-GS framework demonstrates notable advantages and specific application scenarios that clearly distinguish it from existing single-GPU [15, 16, 66, 68, 5], multi-GPU [34, 30], and MVS-based [44] methods. Compared to single-GPU Gaussian Splatting (GS) approaches, ULSR-GS exhibits exceptional scalability and efficiency, making it particularly suitable for large-scale urban reconstruction tasks. The inherent memory constraints of single-GPU methods severely limit their capacity for handling extensive datasets; our multi-GPU strategy circumvents this limitation by employing an optimized point-to-photo partitioning technique driven by SfM. Consequently, ULSR-GS effectively reconstructs detailed urban scenes at scales previously unattainable with single-GPU methods [15, 16, 66, 68, 5].
Furthermore, compared with other multi-GPU GS approaches [34, 30] that typically rely on image-centric partitioning, ULSR-GS provides greater adaptability and accuracy, especially beneficial in complex scenarios such as close-range and oblique photogrammetry. Our innovative per-point partitioning ensures that each sub-region is reconstructed using its most relevant image set, thus addressing critical issues of uneven coverage and irregular camera paths encountered by traditional methods. These improvements are validated by quantitative metrics and qualitative evaluations across diverse urban scenes.
Moreover, relative to MVS-based methods [44], ULSR-GS is particularly effective in reconstructing thin and complex architectural structures, which often present significant challenges for correspondence-based MVS techniques due to matching ambiguities or limited local texture.
Despite these strengths, ULSR-GS also exhibits several limitations: (1) As shown in Figure 10, the framework is less effective in vegetation-rich environments, where dense correspondence-based MVS techniques better capture the intricate and semi-transparent structures inherent in foliage. (2) The generated meshes tend to be overly smooth, potentially lacking essential high-frequency geometric details needed for accurately representing intricate textures and detailed surfaces. (3) The current GS-based training and rendering resolution limit (approximately 1.6K) is insufficient for applications demanding high-detail fidelity, such as high-resolution (2cm per pixel) oblique photogrammetry. (4) Additionally, the existing depth-projection-based densification strategy primarily enhances regions already adequately reconstructed, inadequately addressing areas that are initially sparse or poorly captured.
7 Conclusion
We present ULSR-GS, a framework dedicated to high- fidelity surface extraction in ultra-large-scale scene. Specifically, our partitioning approach combines with a multi-view selection strategy. Additionally, ULSR-GS employs a multi-view geometric consistency densification to enhance surface details. Experimental results demonstrate that ULSR-GS outperforms other SOTA GS-based works on large-scale benchmark datasets. Our approach addresses key limitations of image-based partitioning methods [29, 33, 20] by ensuring finer granularity in mesh boundaries and prioritizing the reconstruction of key scene regions. This strategy enables us to extract meshes individually rather than merge sub-regions and extract the full scene. Unlike existing MVS-based approaches, our method enables the simultaneous generation of a high-quality radiance field for rendering and a detailed mesh representation for reconstruction, all within a single training pipeline. Also, comparison with traditional MVS-based methods demonstrates ULSR-GS drastically reduces processing time while maintaining fine structural details. These advantages make ULSR-GS a promising choice for oblique photogrammetry, offering an efficient and scalable solution for real-world applications.
For future work, we identify three key directions to enhance our framework. First, improving depth computation will be crucial for refining surface extraction and reducing artifacts in complex environments. Second, developing highly efficient 4K-resolution rendering and surface extraction will address the current limitations of GS-based methods in photogrammetric applications. Finally, aerial and terrestrial photogrammetry are combined to complete the detailed reconstruction of the building facade structure while solving the surface reconstruction of the translucent glass building.
8 Acknowledgements
This work was supported by the Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004), the XJTLU AI University Research Centre, the Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU, and the SIP AI Innovation Platform (YZCXPT2022103).